3. You are about to be dealt a 5 card poker hand from a well shuffled deck
A. How many ways can you have a different 5 card hand

Answer by addingup(3677)

1) repetitions are not allowed (e.g. you can't have two ace of diamonds)
2) Order is not important
Then:
n!/[(n-r)!r!]
:
52!/[(52-5)!5]
:
52!/47!5!
:
52*51*50*49*48/5!
:
52*51*(5*5*2)*49*(4*4*3):/5*4*3*2
:
52*51*5*49*4/1 = 52*51*5*49*4 = 2,598,960